a 515kg rollercoaster is at the bottom of a loop with a radius of 10m. If the speed at the bottom of the loop is 20m/s, what is the force of the track pushing up on the vehicle at this point?
1
Expert's answer
2018-12-05T15:28:10-0500

N-mg=ma_c (where N=P)

a_c=v^2/R

N=P=m·(v^2/R-g)=515·(〖20〗^2/10-10)=15450 N